\begin{thebibliography}{33}

\bibitem{alperinbroue}
Alperin J.L. and Brou\'e M., Local methods in block theory,
{\it Ann. of Math\/} {\bf 110} (1979), 143--157.

\bibitem{bcglo1}
Broto C., Castellana N., Grodal J., Levi R. and Oliver B., Subgroup families controlling $p$-local finite groups,
{\it Proc. London Math. Soc. (3)\/} {\bf 91} (2005), no. 2, 325--354.

\bibitem{bcglo}
Broto C., Castellana N., Grodal J., Levi R. and Oliver B., Extensions of $p$-local finite groups,
{\it Trans. Amer. Math. Soc.\/} {\bf 359} (2007), no. 8, 3791--3858.

\bibitem{blo2}
Broto C., Levi R. and Oliver B., Homotopy equivalences of $p$-completed classifying spaces of finite groups,
{\it Invent. Math.\/} {\bf 151} (2003), 611--664.

\bibitem{blo}
Broto C., Levi R. and Oliver B., The homotopy theory of fusion systems,
{\it J. Amer. Math. Soc.\/} {\bf 16} (2003),  no. 4, 779--856.

\bibitem{gl68}
Glauberman G., A characteristic subgroup of a $p$-stable group,
{\it Canad. J. Math.\/} {\bf 20} (1968), 1101--1135.

\bibitem{gl71}
Glauberman G., Global and local properties of finite groups, Finite simple groups,
({\it Proc. Instructional Conf., Oxford, 1969\/}), Academic Press, London, 1971, 1--64.

\bibitem{gl77}
Glauberman G., Factorizations in local subgroups of finite groups,
Amer. Math. Soc., Providence, RI, 1977, Regional Conference Series in Mathematics, No. 33.

\bibitem{gls2b}
Gorenstein D., Lyons R. and Solomon R., The classification of the finite simple groups. Number 2. Part I. Chapter G, Mathematical Surveys and Monographs, vol. 40, Amer. Math. Soc., Providence, RI, 1996, General Group Theory.

\bibitem{kl05}
Kessar R. and Linckelmann M., $ZJ$-theorems for fusion systems,
{\it Trans. Amer. Math. Soc.\/} {\bf 360} (2008), 3093--3106.

\bibitem{klr}
Kessar R., Linckelmann M. and Robinson G.R., Local control in fusion systems of $p$-blocks of finite groups,
{\it J. Algebra\/} {\bf 257} (2002), no. 2, 393--413.

\bibitem{kessarstancu}
Kessar R. and Stancu R., A reduction theorem for fusion systems of blocks,
{\it J. Algebra\/} {\bf 319} (2008), 806--823.

\bibitem{ks04}
Kurzweil H. and Stellmacher B., The theory of finite groups. An introduction,
Springer-Verlag, New York, 2004.

\bibitem{markus}
Linckelmann M., Simple fusion systems and the Solomon $2$-local groups,
{\it J. Algebra\/} {\bf 296} (2006), 385--401.

\bibitem{linckelmann}
Linckelmann M., Fusion Category Algebras,
{\it J. Algebra\/} {\bf 277} (2004), 222--235.

\bibitem{intro}
Linckelmann M., Introduction to fusion systems. Group representation theory,
EPFL Press, Lausanne, 2007, 79--113.

\bibitem{oliver}
Oliver B. Equivalences of classifying spaces completed at odd primes.
{\it Math. Proc. Cambridge Philos. Soc.\/} {\bf 137(2)} (2004), 321–-347.

\bibitem{oliver2}
Oliver B. Equivalences of classifying spaces completed at the prime two.
{\it Mem. Amer. Math. Soc.\/} {\bf 180(848)} (2006), vi+102.

\bibitem{puig:notes}
Puig L., Unpublished notes.

\bibitem{puig:frobenius}
Puig L., Full Frobenius systems and their localizing categories,
{\it Preprint}, June 2001.

\bibitem{puig:fusioncategories}
Puig L., Frobenius Categories,
{\it J. Algebra\/} {\bf 303} (2006), 309-357

\bibitem{ruizviruel}
Ruiz A. and Viruel A., The Classification of $p$-local Finite Groups over the
Extraspecial Group of Order $p^3$ and Exponent $p$, {\it Math. Z.}, {\bf 248} (2004), 45--65.

\bibitem{stancu}
Stancu R., Control of the fusion in fusion systems,
{\it J. Algebra and Applications} {\bf 5} (2006), no.6, 817--837.

\bibitem{st92}
Stellmacher B., An analogue to Glauberman's $ZJ$-theorem,
{\it Proc. Amer. Math. Soc.} {\bf 109} (1990), no. 4, 925-929.

\bibitem{st92b}
Stellmacher B., Errata to: ``An analogue to Glauberman's $ZJ$-theorem''
[{\it Proc. Amer. Math. Soc.\/} {\bf 109} (1990), no.4, 925-929],
Proc. Amer. Math. Soc. {\bf 114} (1992), no. 2, 588.

\bibitem{st96}
Stellmacher B., A characteristic subgroup of $\Sigma\sb 4$-free groups,
{\it Israel J. Math.\/} {\bf 94} (1996), 367-379.

\end{thebibliography}

\documentclass[12pt]{amsart}
\usepackage{amssymb}

\newtheorem{thm}{Theorem}[section]

\newtheorem{lem}[thm]{Lemma}
\newtheorem{prop}[thm]{Proposition}

\theoremstyle{remark}

\newtheorem{rem}[thm]{Remark}
\newtheorem{dfn}[thm]{Definition}
\newtheorem{thma}[thm]{Theorem}
\newtheorem{propa}[thm]{Proposition}
\newtheorem{lema}[thm]{Lemma}

\def\Aut{{\rm Aut}}
\def\Syl{{\rm Syl}}
\def\id{{\rm id}}

\def\Out{{\rm Out}}
\def\Hom{{\rm Hom}}

\def\N#1#2{N_{#1}(#2)}
\def\C#1#2{C_{#1}(#2)}

\def\Cc{{\mathcal C}}

\def\Ff{{\mathcal F}}
\def\Ee{{\mathcal E}}
\def\Gg{{\mathcal G}}
\def\Uu{{\mathcal U}}

\begin{document}

\title[A characteristic subgroup for fusion systems]{A characteristic subgroup for fusion systems}

\author{Silvia Onofrei}
\author{Radu Stancu}
\address{Department of Mathematics, The Ohio State University, 100 Mathematics Tower, 231 West $18^{\rm th}$ Avenue, Columbus, Ohio 43210, USA, Email: onofrei@math.ohio-state.edu}
\address{Department of Mathematical Sciences, University of Copenhagen
Universitetsparken 5, DK-2100 Copenhagen {\O}, Denmark, Email: stancu@math.ku.dk}
\date{\today}
\subjclass{20D15, 20E25}

\begin{abstract}As a counterpart for the prime $2$ to Glauberman's $ZJ$-theorem, Stellmacher proves that any nontrivial $2$-group $S$ has a nontrivial characteristic subgroup $W(S)$ with the following property. For any finite $\Sigma_4$-free group $G$, with $S$ a Sylow $2$-subgroup of $G$ and with $O_2(G)$ self-centralizing, the subgroup $W(S)$ is normal in $G$. We generalize Stellmacher's result to fusion systems. A similar construction of $W(S)$ can be done for odd primes and gives rise to a Glauberman functor.
\end{abstract}

\maketitle

%%%%%%%%%%%%%%%%%%%%%%%%
\section{Introduction}%%
%%%%%%%%%%%%%%%%%%%%%%%%

A fundamental result in the theory of finite groups is Glauberman's $ZJ$-theorem \cite{gl68}. For $p$ be an odd prime, $G$ a finite group and $S$ a Sylow $p$-subgroup of $G$, the $ZJ$-theorem asserts that the center of the Thompson group $Z(J(S))$ is normal in $G$ whenever $G$ is $Qd(p)$-free and $\C G{O_p(G)} \leq O_p(G)$. Recall that for a finite $p$-group $Q$, the Thompson subgroup $J(Q)$ is the subgroup generated by the abelian subgroups of $Q$ of largest order and that the group $Qd(p)= (\mathbb{Z}_p \times \mathbb{Z}_p):SL(2,p)$ is the extension of the $2$-dimensional vector space over $\mathbb{F}_p$ (the field with $p$ elements) by $SL(2,p)$ with its natural action on this vector space. A group $G$ is $H$-free if no section of $G$ is isomorphic to $H$; see also Section~$3$.

More recently, a proof of the $ZJ$-theorem, in the context of fusion systems, was given by Kessar and Linckelman \cite{kl05}. The authors introduce the notion of $Qd(p)$-free fusion system and prove that if $\Ff$ is a $Qd(p)$-free fusion system on a finite $p$-group $S$, with $p$ an odd prime, then $\Ff$ is controlled by $W(S)$, for any Glauberman functor $W$. The related notions of characteristic $p$-functor and Glauberman functor were initially defined in \cite[Definition 1.3]{klr}; they are given below in Definition \ref{charfunct}.

For $p=2$ the $ZJ$-theorem does not hold anymore. In \cite[Question 16.1]{gl71}, Glauberman asks whether there exists a subgroup which is characteristic in a Sylow $2$-subgroup $S$ of a $\Sigma_4$-free group $G$, with the property $\C G{O_p(G)} \leq O_p(G)$. Here $\Sigma_4$ denotes the symmetric group on four letters.

The answer to Glauberman's question was given by Stellmacher \cite{st96}, who also obtained a different proof of the $ZJ$-theorem \cite{st92,st92b}. Stellmacher's idea is to approximate such a subgroup via subgroups of $Z(J(S))$; see \cite[Section 9.4]{ks04} for an overview of this approach. The main theorem in \cite{st96} (see also 6.4 in the Appendix) can be phrased as follows:

\noindent{\it Theorem (Stellmacher): Let $S$ be a finite nontrivial $2$-group. Then there exists a nontrivial characteristic subgroup $W(S)$ of $S$ which is normal in $G$, for every finite $\Sigma_4$-free group $G$ with $S$ a Sylow $2$-subgroup and $\C G{O_2(G)} \leq O_2(G)$.}

Remark that the condition (III) in \cite{st96} is not necessary. A proof of this fact uses Lemmas 6.5 and 6.6 and Remark 6.7 in the Appendix.

In this paper we generalize Stellmacher's approach to fusion systems. Our main result is a proof of Stellmacher's version of the $ZJ$-theorem in the context of fusion systems:

\begin{thm}\label{mainS4}Let $S$ be a finite $2$-group and let $\Ff$ be a $\Sigma_4$-free fusion system over $S$. Then there exists a nontrivial characteristic subgroup $W(S)$ of $S$ with the property that $\Ff=\N \Ff {W(S)}$.
\end{thm}

Since Stellmacher's construction of $W(S)$ gives rise to a Glauberman functor (see Section $4$ for details) we can combine Theorem $B$ in \cite{kl05} with Theorem \ref{mainS4} in our paper to obtain the more general result which is independent of the nature of the prime $p$:

\begin{thm}\label{Qdpfree}Let $S$ be a finite $p$-group and let $\Ff$ be a $Qd(p)$-free fusion system over $S$. Then there exists a nontrivial characteristic subgroup $W(S)$ of $S$ with the property that $\Ff=\N \Ff {W(S)}$.
\end{thm}

Using the same construction for $W(S)$ as in the above theorem, the normal complement theorem due to Thompson \cite[9.4.7]{ks04} can be phrased as:

{\it Theorem (Thompson): Let $G$ be a finite group, $p$ an odd prime and $S$ a Sylow $p$-subgroup of $G$. Then $G$ has a normal $p$-complement provided $\N G{W(S)}$ has such a complement}.

Our third result generalizes Thompson's theorem to the class of fusion systems. This result is similar to Theorem $A$ in \cite{kl05}, except that we replace the group $Z(J)$ with the group $W(S)$ defined in Section~$4$:

\begin{thm}\label{pcomp}Let $\Ff$ be a fusion system over a finite $p$-group $S$, with $p$ an odd prime. Then $\Ff = \Ff_S(S)$ if and only if $\N \Ff {W(S)} = \Ff_S(S)$.
\end{thm}

The paper is organized as follows. Section $2$ contains background material on fusion systems. In Section $3$, the notions of $H$-free fusion system, characteristic $p$-functor and Glauberman functor are defined; further properties of fusion systems are discussed. The characteristic subgroup $W(S)$ is constructed, via two different methods, in Section $4$. The proofs of the theorems are given in Section $5$. In the Appendix a few related results from group theory are included.

%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
\section{Background on Fusion Systems}%%
%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%

Fusion systems were introduced by Puig in 1990 \cite{puig:notes,puig:frobenius} in an effort to axiomatize the $p$-local structure of a finite group and of a block of a group algebra - the work was published only recently \cite{puig:fusioncategories} but was known to the community long before. In 2000 Broto, Levi and Oliver \cite{blo} 
enriched this axiomatic approach by introducing the centric linking system that is essentially linked to the associated $p$-completed classifying space. The third author used this axiomatic frame to reformulate and solve the Martino-Priddy conjecture \cite{oliver,oliver2}. Broto, Levi and Oliver give a different definition of the fusion systems which they proved to be equivalent to Puig's definition. In this paper we use a simplified definition which we find more elegant, equivalent to the above ones \cite{kessarstancu}.

We start with a more general definition, following \cite{linckelmann}.

A {\it category $\Ff$ on a finite $p$-group $S$} is a category whose objects are
the subgroups of $S$ and whose set of morphisms
between the subgroups $Q$ and $R$ of $S$, is a set $\Hom_\Ff(Q,R)$ of
injective group homomorphisms from $Q$ to $R$, with the following properties:
\begin{enumerate}
\item[(1)] if $Q\le R$ then the inclusion of $Q$ in $R$ is a morphism in $\Hom_\Ff(Q,R)$;
\item[(2)] for any $\varphi\in\Hom_\Ff(Q,R)$ the induced isomorphism $Q\simeq\varphi(Q)$
and its inverse are morphisms in $\Ff$;
\item[(3)] the composition of morphisms in $\Ff$ is the
usual composition of group homomorphisms.
\end{enumerate}

Let $\Ff_1$ be a category on $S_1$ and $\Ff_2$ a category on $S_2$.
A {\it morphism} between $\Ff_1$ and $\Ff_2$ is a pair $(\alpha, \Theta)$ with $\alpha \in \Aut(S)$ and
$\Theta: \Ff_1 \rightarrow \Ff_2$ a covariant functor, such that:
\begin{itemize}
\item[(i)]for any subgroup $Q$ of $S$, $\alpha (Q) = \Theta(Q)$;
\item[(ii)]for any morphism $\varphi$ in $\Ff_1$, $\Theta(\varphi)\circ \alpha = \alpha \circ \varphi$.
\end{itemize}

In the following we give a series of useful definitions in a category
$\Ff$ on $S$. If there exists an isomorphism $\varphi\in\Hom_\Ff(Q,R)$ we say that
$Q$ and $R$ are {\it $\Ff$-conjugate}.

We say that a subgroup $Q$ of $S$ is
\begin{list}{\upshape\bfseries}
{\setlength{\leftmargin}{.8cm}
\setlength{\rightmargin}{0cm}
\setlength{\labelwidth}{0.8cm}
\setlength{\labelsep}{0.2cm}
\setlength{\parsep}{0.5ex plus 0.1ex minus 0.1ex}
\setlength{\itemsep}{0.2ex plus 0.1ex minus 0.1ex}}
\item[(i)] {\it fully $\Ff$-centralized} if $|\C SQ|\ge|\C S{Q'}|$ for all $Q'\le S$
which are $\Ff$-conjugate to $Q$.
\item[(ii)] {\it fully $\Ff$-normalized} if $|\N SQ|\ge|\N S{Q'}|$ for all $Q'\le S$
which are $\Ff$-conjugate to $Q$.
\item[(iii)] {\it $\Ff$-centric} if $\C S{\varphi(Q)}\subseteq \varphi(Q)$, for all $\varphi\in\Hom_\Ff(Q,S)$.
\item[(iv)] {\it $\Ff$-radical} if $O_p(\Out_\Ff(Q))=1$.
\item[(v)] {\it $\Ff$-essential} if $Q$ is $\Ff$-centric and $\Out_\Ff(Q)$
has a strongly $p$-embedded proper subgroup $M$
(that is $M$ contains a Sylow $p$-subgroup $P$ of $\Out_\Ff(Q)$ such that
$P\ne\,^\varphi P$ and $^\varphi P\cap P=\{1\}$ for every $\varphi\in\,\Out_\Ff(Q)\setminus M$ ).
\end{list}

For $Q,\,R \leq S$ we denote $\Hom_S(Q,R):=\{u\in S\,|\,^uQ\le R\}/\C SQ$
and $\Aut_S(Q):=\Hom_S(Q,Q)$. Other useful notations are $\Aut_\Ff(Q):=\Hom_\Ff(Q,Q)$
and $\Out_\Ff(Q):=\Aut_\Ff(Q)/\Aut_Q(Q)$.

We are now ready to give the definition of a fusion system.

A {\it fusion system} on a finite $p$-group $S$ is a category $\Ff$ on $S$ satisfying the following
properties:
\begin{enumerate}
\item[FS1.] $\Hom_S(Q,R)\subseteq\Hom_\Ff(Q,R) \text{ for all } Q,R\le S\,.$

\item[FS2.] $\Aut_S(S)$ is a Sylow $p$-subgroup of $\Aut_\Ff(S)$.

\item[FS3.] Every $\varphi:Q\to S$ such that $\varphi(Q)$ is fully $\Ff$-normalized extends to a morphism $\widehat\varphi:N_\varphi\to S$ where
$$N_\varphi=\{x\in\N SQ\,|\,\exists y\in\N S{\varphi(Q)},\,\varphi(\,^xu)=\,^y\varphi(u),\,\forall u\in Q\}\,.$$
\end{enumerate}

Remark that $N_\varphi$ is the largest subgroup of $\N SQ$ such that
$^\varphi(N_\varphi/\C SQ)\le\Aut_S(\varphi(Q))$. Thus we always have $Q\C SQ\le N_\varphi \le N_S(Q)$.

If $\Ff$ is a fusion system on $S$ and $Q\le S$ we have the following
equivalent characterization of being fully $\Ff$-normalized.

\begin{prop}[\cite{markus}, Proposition 1.6]\label{prpsylow}
A subgroup $Q$ of $S$ is fully $\Ff$-normalized if and only if $Q$ is fully
$\Ff$-centralized and $\Aut_S(Q)$ is a Sylow $p$-subgroup of $\Aut_\Ff(Q)$.
\end{prop}

In the following Lemma we recollect two useful properties involving fully $\Ff$-normalized subgroups; see also \cite[Lemmas 2.2, 2.3]{kl05}. For completeness we include the proofs.

\begin{lem}\label{fullynormalized} Let $\Ff$ be a fusion system on a finite $p$-group $S$ and $Q$ a subgroup of $S$.
\begin{itemize}
\item[a)] There is a morphism $\varphi \in \Hom_\Ff(N_S(Q),S)$ such that $\varphi(Q)$ is fully $\Ff$-normalized.
\item[b)] If $Q$ is fully $\Ff$-normalized, then $\varphi(Q)$ is fully normalized, for any morphism $\varphi \in \Hom_\Ff(N_S(Q),S)$.
\end{itemize}
\end{lem}

\begin{proof}
a) Let $\psi:Q \rightarrow S$ be a morphism with $\psi(Q)$ fully $\Ff$-normalized. By Proposition $2.1$, $\Aut_S(\psi(Q))$ is a Sylow $p$-subgroup of $\Aut_\Ff (\psi(Q))$. Since $\psi \circ \Aut_S(Q) \circ \psi^{-1}$ is a $p$-subgroup of $\Aut_\Ff(\psi(Q))$ it follows that there exists a morphism $\tau \in \Aut_\Ff(\psi(Q))$ with $\tau \psi \circ \Aut_S(Q) \circ \psi^{-1} \tau^{-1} \le \Aut_S(\psi(Q))$. Set $\alpha = \tau \psi$ and observe that $\alpha(Q)$ is fully $\Ff$-normalized. By the extension axiom FS3, $\alpha$ extends to a morphism $\varphi:N_\alpha \rightarrow S$. But since $\alpha \circ \Aut_S(Q) \circ \alpha^{-1} \le \Aut_S(\psi(Q))$ it follows that $N_{\alpha} = N_S(Q)$. Henceforth there exists a morphism $\varphi \in \Hom _\Ff (\N S Q, S)$ such that $\varphi(Q)$ is fully $\Ff$-normalized.\\
b) Since $Q$ is fully $\Ff$-normalized and since $\varphi$ is a morphism in $\Ff$, hence injective, we have: $\varphi(N_S(Q))=N_S(\varphi(Q))$.
\end{proof}

Puig \cite{puig:notes} gave analogous notions for the normalizer and the centralizer in a finite group:

The {\it normalizer of $Q$ in $\Ff$} is the category $\N\Ff Q$ on $\N SQ$
having as morphisms those morphisms $\varphi\in\Hom_{\Ff}(R,T)$, for $R$ and $T$ subgroups of $\N SQ$, satisfying that there exists a morphism $\widehat{\varphi}\in\Hom_\Ff(QR,QT)$ such that $\widehat{\varphi}|_Q\in\Aut_\Ff(Q)$ and $\widehat{\varphi}|_R=\varphi$. If $Q \le S$ has the property that $\Ff=\N\Ff Q$ then we say that $Q$ {\it is normal in} $\Ff$. The largest subgroup of $S$ which is normal in $\Ff$ will be denoted $O_p(\Ff)$.

The {\it centralizer of $Q$ in $\Ff$} is the category $\C\Ff Q$ on $\C SQ$
having as morphisms those morphisms $\varphi\in\Hom_\Ff(R,T)$, with $R$ and $T$ subgroups of $C_S(Q)$, satisfying that there exists a morphism $\widehat{\varphi}\in\Hom_\Ff(QR,QT)$ such that $\widehat{\varphi}|_Q=\id_Q$ and $\widehat{\varphi}|_R=\varphi$.

Also denote by $\N SQ\C\Ff Q$ the category on $\N SQ$ having as morphisms all group homomorphisms $\varphi: P \rightarrow R$, for $P$ and $R$ subgroups of $\N SQ$, for which there exists a morphism $\psi: QP \rightarrow QR$ and $x \in\N SQ$ such that $\psi|_Q = c_x$ (the morphism induced by conjugation by $x$) and $\psi|_P=\varphi$.

\begin{prop}[\cite{puig:frobenius}, Proposition 2.8]\label{prpnorm}
If $Q$ is fully $\Ff$-normalized then $\N\Ff Q$ is a fusion system on $\N SQ$. 
If $Q$ is fully $\Ff$-centralized then $\C\Ff Q$ and $\N SQ\C \Ff Q$ are fusion systems on $\C SQ$.
\end{prop}

Alperin's theorem on $p$-local control of fusion also holds for fusion systems. First we set up this theorem's notations and terminology. If $\varphi\in\Aut_\Ff(S)$ we say that $\varphi$ is a maximal $\Ff$-automorphism. If $\varphi\in\Aut_\Ff(E)$, with $E$ an $\Ff$-essential subgroup of $S$, we say that $\varphi$ is an essential $\Ff$-automorphism. Alperin's fusion theorem asserts that the essential and maximal $\Ff$-automorphisms suffice to determine the whole fusion system $\Ff$.

\begin{thm}[Alperin]\label{AlperinsThm} Any morphism $\varphi\in\Hom_\Ff(Q,S)$ can be written as the composition of restrictions of essential $\Ff$-automorphisms, followed by the restriction of a maximal $\Ff$-automorphism.
More precisely, there exists
\begin{itemize}
\item[(a)] an integer $n\ge 0$,
\item[(b)] a set of $\Ff$-isomorphic subgroups of $S$, $Q=Q_0,\,Q_1,\,\dots,\,Q_n,\,Q_{n+1}=\varphi(Q)$,
\item[(c)] a set of $\Ff$-essential, fully $\Ff$-normalized subgroups
$E_i$ of $S$ containing $Q_{i-1}$ and $Q_i$, for all $1\le i\le n$,
\item[(d)] a set of essential automorphisms $\psi_i\in\Aut_\Ff(E_i)$ satisfying $\psi_i(Q_{i-1})=Q_i$, for all $1\le i\le n$ and
\item[(e)] a maximal automorphism $\psi_{n+1}\in\Aut_\Ff(S)$ satisfying $\psi_i(Q_n)=Q_{n+1}$,
\end{itemize}
such that we have
$$\varphi(u)=\psi_{n+1}\psi_{n}\ldots\psi_1(u),\text{ for all }u\in Q\,.$$
\end{thm}

The reader can find a proof of this theorem in \cite{stancu}; an alternative proof of this theorem in a different axiomatic setting was given by Puig \cite[Corollary 3.9]{puig:frobenius} and another in a less general form, using $\Ff$-centric, $\Ff$-radical subgroups instead of $\Ff$-essential subgroups, can be found in \cite[Theorem A.10]{blo}. We use this later form in the proof of Lemma \ref{Op_cr}\,. 

The classical examples of a fusion systems are the ones coming from the $p$-local structure of a finite group $G$. If $S$ is a Sylow $p$-subgroup of $G$ then we denote by $\Ff_S(G)$ the fusion system on $S$ having as morphisms
$$\Hom_{\Ff_S(G)}(P,Q)=\N G{P,Q}/\C GP =\Hom_G(P,Q)$$
where $P,\,Q$ are subgroups of $S$ and $\N G{P,Q}=\{g\in G\,|\,^gP\le Q\}$ is the $G$-transporter from $P$ to $Q$.

There are examples of fusion systems that do not come from a finite group (see eg. \cite{ruizviruel}). But there are also particular cases when one can construct a finite group with $p$-local structure equivalent to a given fusion system. This is the case for constrained fusion systems. The fusion system $\Ff$ is said to be {\it constrained} if $O_p(\Ff)$ is $\Ff$-centric. Any constrained fusion system was proven to come from a finite group by Broto, Castellana, Grodal, Levi and Oliver:

\begin{thm}\cite[Theorem 4.3]{bcglo1}\label{thmconstraint}
Let $\Ff$ be a fusion system on $S$ and suppose that there exists an $\Ff$-centric subgroup $Q$ of $S$
such that $\N\Ff Q=\Ff$ (in particular $\Ff$ is constrained). Then there exists a, unique up to isomorphism,
finite $p'$-reduced $p$-constrained group $L_Q^{\Ff}$ (i.e $O_{p'}(L_Q^{\Ff})=1$ and $Q$ is a normal 
subgroup of $L_Q^{\Ff}$) having $S$ as a Sylow $p$-subgroup and such that
$\Ff= \Ff_S (L_Q^{\Ff})$. Furthermore $L_Q^{\Ff}/Z(Q)\simeq\Aut_{\Ff}(Q)$.
\end{thm}

%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
\section{Further results on fusion systems} %%
%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%

Let $G$ be a finite group and $p$ a prime divisor of its order. If $A \trianglelefteq B \leq G$ then $B/A$ is called a {\it section} of $G$. We say that $H$ {\it is involved in} $G$ if $H$ is isomorphic to a section of $G$. If $H$ is not involved in $G$ then $G$ is called $H$-{\it free}.

Following \cite{kl05} we say that the {\it fusion system $\Ff$ on $S$ is $H$-free} if $H$ is not involved in any of the groups $L_Q^{\Ff}$, for $Q$ running over the set of 
$\Ff$-centric, $\Ff$-radical and fully $\Ff$-normalized subgroups of $S$. In some particular cases the property of being $H$-free passes to subsystems and quotient systems as the next two results from \cite{kl05} show.

\begin{prop}\cite[Proposition 6.3]{kl05}\label{subsyst} Let $\Ff$ be a fusion system on a finite $p$-group $S$ and let $Q$ be a fully $\Ff$-normalized subgroup of $S$. If $\Ff$ is $H$-free, then so is any fusion subsystem of $\Ff$ which lies between $\N\Ff Q$ and $\N SQ\C\Ff Q$. In particular, if $\Ff$ is $H$-free, so are $\N \Ff Q$ and $\N SQ \C\Ff Q$.
\end{prop}

Let $\Ff$ be a fusion system on $S$ and let $Q$ be a subgroup of $S$ with the property $\Ff = \N \Ff Q$. The category $\Ff /Q$ on $S/Q$ is defined as follows: for $Q \leq P,R \leq S$, a group homomorphism $\psi: P/Q \rightarrow R/Q$ is a morphism in $\Ff/Q$ if there is a morphism $\varphi \in \Hom_\Ff(P,R)$ satisfying $\psi(xQ) = \varphi(x)Q$ for all $x \in P$. The fact that $\Ff/Q$ is a fusion system on $S/Q$ is due to Puig \cite{puig:notes}, see also \cite[Proposition 2.8]{kl05}.

\begin{prop}\cite[Proposition 6.4]{kl05}\label{quotient} Let $\Ff$ be a fusion system on a finite $p$-group $S$ and let $Q$ be a normal subgroup of $S$ such that $\Ff=\N \Ff Q$. If $\Ff$ is $H$-free then $\Ff /Q$ is also $H$-free.
\end{prop}

The following result generalizes the main technical step in the proof of \cite[Proposition 5.2]{kl05}.

\begin{prop}\label{wellplaced} Let $\Ff$ be a fusion system on a finite $p$-group $S$ and let $W_i$,
$1 \leq i \leq n$ be subgroups of $S$ such that
\begin{itemize}
\item[(a)] the subgroup $W_{i+1}$ is a characteristic subgroup of $\N S{W_i}$ for $1 \leq i \leq n-1$;
\item[(b)] the subgroup $W_i$ is fully $\Ff$-normalized for $1 \leq i \leq n-1$.
\end{itemize}
Then there exists a morphism $\varphi \in \Hom _\Ff (\N S{W_n}, S)$ such that $\varphi(W_i)$ is fully $\Ff$-normalized for all $1 \leq i \leq n$. In particular $\varphi(\N S{W_i})=\N S{\varphi (W_i)}$. If, moreover, $W_i$ is $\Ff$-centric and/or $\Ff$-radical for some $1 \le i \le n$, then so is $\varphi(W_i)$.
\end{prop}

\begin{proof}
If follows from Lemma $2.2(a)$ that there exists a morphism $\varphi\in\Hom _{\Ff} (N_S(W_n), S)$ such that $\varphi(W_n)$ is fully $\Ff$-normalized. According to condition (a), $W_{i+1}$ is a characteristic subgroup of $N_S(W_i)$, for $1 \leq i \le n-1$ and therefore $N_S(W_i) \le N_S(W_{i+1})$. In particular $N_S(W_i) \leq N_S(W_n)$ and the morphism $\varphi$ is defined on $\N S{W_i}$ for all $1 \leq i \leq n$.

Next we show $\varphi(W_i)$ is fully $\Ff$-normalized for all $1 \le i \le n-1$. By elementary group theory $\varphi(N_S(W_i)) \le N_S(\varphi(W_i))$. Since $\varphi$ is injective it follows that $|N_S(W_i)| \le |\varphi(N_S(W_i))|$. But, according to (b), $W_i$ is fully $\Ff$-normalized and $|N_S(W_i)| \ge |N_S(\varphi(W_i))|$. It follows now that $|N_S(W_i)| = |N_S(\varphi(W_i))|$ which shows that $\varphi(W_i)$ is fully $\Ff$-normalized and that $\varphi (N_S(W_i)) = N_S(\varphi(W_i))$, for all $1 \le i \le n-1$.

Since $W_i$ is fully $\Ff$-normalized for all $1 \leq i \leq n$, it is also fully $\Ff$-centralized. Thus $\varphi (\C S{W_i})=\C S{\varphi (W_i)}$ and if $W_i$ is $\Ff$-centric, then so is $\varphi(W_i)$. Moreover it is a general fact that $\,^\varphi\Aut_\Ff(W_i)=\Aut_\Ff(\varphi(W_i))$
and $\,^\varphi\Aut_S(W_i)=\Aut_S(\varphi(W_i))$ so if $W_i$ is $\Ff$-radical, then so is $\varphi(W_i)$.
\end{proof}

We need the following definition reproduced from \cite[Definition 5.1]{kl05}.

\begin{dfn}\label{charfunct}
A {\it positive characteristic functor} is a map sending any nontrivial finite $p$-group $S$ to a nontrivial characteristic subgroup $W(S)$ of $S$ such that $W(\varphi(S)) = \varphi(W(S))$ for every $\varphi \in \Aut(S)$. A positive characteristic functor is a {\it Glauberman functor} if whenever $S$ is a Sylow $p$-subgroup of a $Qd(p)$-free finite group $L$ which satisfies $C_L(O_p(L)) = Z(O_p(L))$, then $W(S)$ is normal in $L$.
\end{dfn}

Using Proposition \ref{wellplaced} we can give a different proof for Proposition $5.3$ in \cite{kl05}.

\begin{prop}\cite[Proposition 5.3]{kl05}\label{normal} Let $\Ff$ be a fusion system on a finite $p$-group $S$ and let $W$ be a positive characteristic functor. Assume that for any non-trivial, proper, $\Ff$-centric, $\Ff$-radical, fully $\Ff$-normalized subgroup $Q$ of $S$ the following holds $\N \Ff Q = \N {\N \Ff Q} {W(\N SQ)}$. Then $\Ff = \N \Ff {W(S)}$.
\end{prop}

\begin{proof}
Suppose that the conclusion does not hold. By Alperin's Fusion Theorem there exists a proper, $\Ff$-centric, $\Ff$-radical, fully $\Ff$-normalized subgroup $Q$ of $S$ such that $\Aut_{\N \Ff {W(S)}}(Q) \subsetneq \Aut_\Ff(Q)$.

Set $W_1=Q$ and define recursively $W_{i+1}:=W(\N S{W_i})$. So $W_{i+1}$ is characteristic in $\N S{W_i}$ implying that we get the following inclusions $\N S{W_i}< \N S{\N S {W_i}} \leq \N S{W_{i+1}}$ this in its turn implies there exists $n\ge 1$ such that the sequence of $\N S{W_i}$ for $1\le i\le n$ is strictly increasing and $\N S{W_n}=S$. Observe that if $\N S{W_i} = \N S{W_{i+1}}$ then $\N S{W_i} = \N S{\N S {W_i}} \leq S$ and by an elementary property of $p$-groups it follows that $N_S(W_i)=S$, thus indeed the sequence eventually reaches $S$.

Moreover the sequence $\{W_i,\,1\le i\le n+1\}$ can be chosen so that all its terms are fully $\Ff$-normalized. This can be done recursively by applying Proposition \ref{wellplaced}, for all $2\le k\le n+1$ to the partial subsequences $\{W_i,\,1\le i\le k\}$. The $W_i$'s are successively modified by replacing them with their images through the morphism $\varphi$ given by Proposition \ref{wellplaced}.

Consider the sequence of the normalizers in $\Ff$ of the $W_i$'s for $1\le i\le n+1$. Given that $W_i$, $1\le i\le n$ are fully $\Ff$-normalized, we have that $\N\Ff{W_i}$ is a fusion system on $\N S{W_i}$. It follows from our assumption that $\N\Ff{W_i}\subseteq\N\Ff{W_{i+1}}$ for all $1 \le i \le n-1$. But then $\N\Ff{Q}=\N\Ff{W_1}\subseteq\N\Ff{W_{n+1}}=\N\Ff{W(S)}$. At the level of morphisms on $Q$ this gives $\Aut_\Ff(Q)\subseteq\Aut_{\N\Ff{W(S)}}(Q)$ which is a contradiction with the initial supposition on $Q$.
\end{proof}

Proposition \ref{normal} is used in \cite{kl05} to prove the next important result.

\begin{prop}\cite[Proposition 3.4]{kl05}\label{quot} Let $S$ be a finite $p$-group and let $Q$ be a normal subgroup of $S$. Let $\Ff$ and $\Gg$ be fusion systems on $S$ such that $\Ff=S \C \Ff Q$ and such that $\Gg \subseteq \Ff$. Let $P$ be a normal subgroup of $S$ containing $Q$. We have $\Gg = \N \Ff P$ if and only if $\Gg /Q = \N {\Ff /Q} {P / Q}$.
\end{prop}

Next we give an application of the Frattini argument to fusion systems. The group theoretic result states that, if $G$ is a finite group, then thefactorization $G=\C GQ\N GR$ holds with $Q=O_p(G)$ and $R=\C G{Q\C SQ}$. This is easily seen to be true as $\C GQ \trianglelefteq G$ and $\N G{\C SQ} \leq \N GR$, then an application of the Frattini argument gives the result.

\begin{lem}\label{Op_cr} Let $\Ff$ be a fusion system on $S$, $Q=O_p(\Ff)$ and $R=Q\C SQ$. Set $\Ff_1=S\C \Ff Q$ and $\Ff_2=\N \Ff R$. Then $\Ff=\langle \Ff_1, \Ff_2 \rangle$.
\end{lem}

\begin{proof}
First remark that $\Ff$, $\Ff_1$ and $\Ff_2$ are fusion systems on $S$ with $\Ff$ containing the other two. By Alperin's fusion theorem (see Theorem \ref{AlperinsThm}), it is enough to prove that for every $\Ff$-centric, $\Ff$-radical, fully $\Ff$-normalized subgroup $U$ of $S$ we have $\Aut_\Ff(U)=\Aut_{\Gg}(U)$ with $\Gg= \langle \Ff_1, \Ff_2 \rangle$.

We shall prove that every $\varphi\in\Aut_\Ff(U)$ can be written as composition of morphisms in $\Ff_1$ and $\Ff_2$ and thus will be contained in $\Aut_{\Gg}(U)$. This will finish our proof as the opposite inclusion is clearly satisfied.

Given that $U$ is $\Ff$-centric and $\Ff$-radical we have by \cite[Proposition 5.6]{intro} that $Q\le U$. Hence $\varphi$ restricts to an automorphism $\theta\in\Aut_\Ff(Q)$. Now we have that $N_\theta$ contains $U$ and $R$ so it contains $UR$. Given that $Q$ is fully $\Ff$-normalized $\theta$ extends to $\chi\in\Hom_\Ff(UR,S)$. Moreover $\chi(R)=\chi(Q)\C S{\chi(Q)}=R$ so in fact $\chi\in\Hom_{\Ff_2}(UR,S)$.

Denote by $\psi$ the restriction to $U$ of $\chi$; then $\psi\in\Hom_{\Ff_2}(U,\psi(U))$.
Both $\varphi$ and $\psi$ restrict as $\theta$ on $Q$ so $\varphi\circ\psi^{-1}$ belongs to $\Hom_{\Ff_1}(\psi(U),U)$. The conclusion in the lemma follows as $\varphi=\varphi\circ\psi^{-1}\circ\psi\in\Aut_{\Gg}(Q)$.
\end{proof}

We close this section with a straightforward result on fusion control in fusion systems.

\begin{lem}\label{normalizerW} Let $W$ be a fully $\Ff$-normalized subgroup of $S$ and suppose that there are two fusion subsystems $\Ff_1$ and $\Ff_2$ of $\Ff$ such that $\Ff=\langle \Ff_1,\Ff_2 \rangle$. If moreover $\Ff_1 = \N {\Ff_1} W$ and $\Ff_2 = \N {\Ff_2} W$. Then $\Ff = \N \Ff W$.
\end{lem}

\begin{proof} We have that $\langle \N {\Ff_1}W, \N {\Ff_2}W \rangle \subseteq \N \Ff W \subseteq \Ff$. The result follows. \end{proof}

%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
\section{A characteristic subgroup of $S$}\label{charsg}%%
%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%

Let $S$ be a finite $p$-group. In this section we construct a subgroup $W(S)$ which is characteristic in $S$ and such that $\Omega(Z(S)) \leq W(S) \leq \Omega(Z(J(S)))$, and with the property that $W(S) \trianglelefteq \Ee$ for all $(\varphi, \Ee) \in \Uu_J$, with $\Uu_J$ a class of embeddings defined below. The notation $\Omega (H)$ stands for the subgroup of $H$ generated by all the elements of order $p$, while $J(S)$ denotes the Thompson subgroup of $S$ defined in Section 1. We shall give below two different, although equivalent, constructions of this characteristic subgroup of $S$ which we shall denote $W(S)$ and $W$. The first construction follows the approach developed by Stellmacher \cite{st92,st96} for finite groups, in which such a subgroup is approximated from various subgroups of $Z(J(S))$. The second construction uses basic properties of fusion systems.

\subsection{The group W(S)} An {\it embedding} is a pair $(\varphi, \Ee)$ where $\varphi \in \Aut(S)$ and $\Ee$ is a category on $\varphi(S)= S$. Let $\Cc$ denote the family of all embeddings of $S$. A nonempty subclass $\Uu$ of $\Cc$ is {\it characteristically closed} if $(\varphi \alpha, \Ee) \in \Uu$ whenever $(\varphi, \Ee) \in \Uu$ and $\alpha \in \Aut(S)$.

An {\it equivalence} between two embeddings $(\varphi _1, \Ee_1)$ and $(\varphi _2, \Ee_2)$ is a morphism
$(\alpha, \Theta):\Ee_1 \rightarrow \Ee_2$ with $\alpha \varphi _1 = \varphi _2$ and
$\Hom_{\Ee_2}(\varphi _2(Q), \varphi _2(R)) = \alpha \circ \Hom _{\Ee_1}(\varphi _1(Q), \varphi_1(R)) \circ {\alpha ^{-1}}_{|\varphi_1 (Q)}$.
The equivalence of embeddings defines an equivalence relation on $\Uu$. Since $S$ is a finite group,
the collection of equivalence classes $[\Uu]$ is a finite set.

Let $O_S(\Uu)$ denote the largest subgroup of $S$ which satisfies the property that $\varphi (O_S(\Uu))$ is normal in $\Ee$ for every embedding $(\varphi, \Ee)$ in $\Uu$.

\begin{lem}Let $\Uu$ be a characteristically closed subclass of $\Cc$ and let $\alpha \in \Aut(S)$. Let $Q$ be a subgroup of $S$ with the property that $\varphi (Q) \trianglelefteq \Ee$ for every $(\varphi, \Ee) \in \Uu$. Then $\varphi(\alpha (Q)) \trianglelefteq \Ee$ for every $(\varphi, \Ee)$. In particular $O_S(\Uu)$ is a characteristic subgroup of $S$.
\end{lem}
\begin{proof} Observe that since $\Uu$ is characteristically closed and $(\varphi, \Ee) \in \Uu$ then $(\varphi\alpha, \Ee) \in \Uu$ and thus $\varphi (\alpha (Q)) \trianglelefteq \Ee$ for every $(\varphi, \Ee) \in \Uu$. The fact that $O_S(\Uu)$ is $\Aut(S)$-invariant follows from its definition.
\end{proof}

Let $\Uu_J$ denote the class of embeddings $(\varphi, \Ee)$ which satisfy the following conditions:
\begin{itemize}
\item[$(U1)$]$\Uu_J$ is characteristically closed.
\item[$(U2)$]$J(\varphi(S))=J(S)$ is normal in $\Ee$ for all $(\varphi, \Ee) \in \Uu_J$.
\item[$(U3)$]$\Ee$ is a $Qd(p)$-free fusion system.
\end{itemize}

For a $p$-group $P$ we set $A(P)=\Omega(Z(P))$ and $B(P)=\Omega(Z(J(P)))$. Remark that $A(P)\le B(P)$ as $Z(P)\le J(P)$. Note that $\alpha(A(P))=A(\alpha(P))$ and $\alpha(B(P))=B(\alpha(P))$ for all $\alpha \in \Aut(P)$, as $A(P)$ and $B(P)$ are characteristic subgroups of $P$.

Define recursively a subgroup $W(S) \leq B(S)$ as follows. Let
$$W_0:= A(S)=\Omega(Z(S)) \leq B(S)$$
and assume that for $i \geq 1$ the subgroups $W_0, W_1, \ldots W_{i-1}$ with
$$W_0 < W_1 <\ldots W_{i-1} \leq B(S)$$
are defined. If $\varphi(W_{i-1})\trianglelefteq \Ee$ for all $(\varphi, \Ee) \in \Uu_J$ then set $W(S):=W_{i-1}$. Otherwise, choose $(\varphi _i, \Ee_i) \in \Uu_J$ to be such that $\varphi_i(W_{i-1})$ is not normal in $\Ee_i$ and define
$$W_i:= \varphi^{-1}_i \left< \varphi_i (W_{i-1})^{\Ee_i} \right> = \varphi^{-1}_i \left< \psi(\varphi _i(W_{i-1})) | \psi \in \Hom_{\Ee_i}(\varphi_i(W_{i-1}),\varphi_i(S)) \right>$$
to be the preimage in $S$ of the group generated by the $\Ee_i$-orbit of $\varphi_i(W_{i-1})$.

Since $B(\varphi_i(S))$ is a characteristic subgroup of $J(\varphi_i(S))$, which is itself normal in $\Ee_i$, it follows that $B(\varphi_i(S))$ is also normal in $\Ee_i$. Clearly $\varphi_i(W_i) \leq B(\varphi_i(S))$ since $\varphi_i(W_i)$ is generated by various conjugates of $\varphi_i(W_{i-1})$ and $\varphi_i(W_{i-1}) \leq B(\varphi_i(S))$ by construction. Thus we have:
$$A(\varphi_i(S)) \leq \varphi_i(W_{i-1}) < \varphi_i(W_i) \leq B(\varphi_i(S)) \trianglelefteq \Ee_i$$ as $A(\varphi_i(S))=\varphi_i(A(S))$ and $B(\varphi_i(S))=\varphi_i(B(S))$ for $\varphi_i \in \Aut (S)$. Then it follows:
$$A(S) \leq W_{i-1} < W_i \leq B(S)$$
As $S$ is finite, this recursive definition terminates after a finite number $n$ of steps and $W(S) := W_n$. Therefore we obtain a chain of subgroups of $Z(J(S))$:
$$A(S) = W_0 < W_1 \ldots < W_i < \ldots < W_n=W(S) \leq B(S)$$
and $\varphi(W(S)) \trianglelefteq \Ee$ for all $(\varphi, \Ee) \in \Uu_J$.

The group $W(S)$ depends on $S$ only and it is independent of the pairs $(\varphi_i, \Ee_i)$. To see this assume that we defined in an analogous way
$$W_0 = \overline{W}_0<\overline{W}_1 < \ldots < \overline{W}_{\tilde{n}} =: \overline{W}(S)$$
for suitable pairs $(\overline{\varphi}_j, \overline{\Ee}_j)$ in $\Uu_J$ and for $j =1, \ldots \overline{n}$. First note that $\overline{W}_0=W_0 \leq W(S) \cap \overline{W}(S)$. Thus $\overline{\varphi}_1(\overline{W}_0)=\overline{\varphi}_1(W_0) \leq \overline{\varphi}_1(W(S)) = W(S)$ as $W(S) \trianglelefteq \overline{\Ee}_1$, and therefore $\overline{\varphi}_1 (\overline{W}_1) = \langle \overline{\varphi}_1 (\overline{W}_0) ^{\overline{\Ee}_1} \rangle \leq W(S)$ which implies $\overline{W}_1 \leq W(S)$. Proceed by induction on $j$; a similar argument shows that since $\overline{W}_{j-1} \leq W(S)$ then $\overline{\varphi}_j(\overline{W}_j) \leq W(S)$ and $\overline{W}_j \leq W(S)$. Therefore $\overline{W}(S) \leq W(S)$. Similarly $W(S) \leq \overline{W}(S)$ and thus $W(S) = \overline{W}(S)$.

\begin{lem}Let $\alpha \in \Aut(S)$. Then $W(\alpha(S)) = \alpha(W(S))$. In particular, $W(S)$ is a characteristic subgroup of $S$, nontrivial if $S$ is nontrivial.
\end{lem}

\begin{proof} The mapping $(\varphi, \Ee) \rightarrow (\varphi \alpha, \Ee)$ defines a bijection on $\Uu_J$. Under this map, the chain of subgroups:
$$A(S) = W_0 < \ldots < W_i < \ldots < W_n = W(S) \leq B(S)$$
is taken to the following chain:
$$A(S) = \alpha(W_0) < \alpha(W_1) < \ldots < \alpha(W_i) < \ldots <\alpha(W_n) = \alpha(W(S)) \leq B(S).$$
Therefore $W(\alpha(S))=\alpha(W(S))$. The last statement follows from the fact that $\Omega(Z(S))\leq W(S)$ and $Z(S) \ne 1$ if $S \ne 1$.
\end{proof}

\subsection{The group W} Denote by $\Cc_J$ the class of categories $\Ff$ on $S$ which satisfy the following conditions:
\begin{itemize}
\item[$(C1)$]$J(S)$ is normal in $\Ff$ for all $\Ff \in \Cc_J$.
\item[$(C2)$]$\Ff$ is a $Qd(p)$-free fusion system.
\end{itemize}

\begin{prop}Let $W_0=\Omega (Z(S))$ and define
$$W:=\langle \psi(W_0) \; | \; \psi \in \Hom_\Ff(J(S), S) \;{\rm for}\;\Ff \in \Cc_J \rangle.$$
The subgroup $W$ is a nontrivial characteristic subgroup of $S$.
\end{prop}

\begin{proof}For all $\alpha \in \Aut(S)$, we will show that $\alpha(W) = W$. Let $\Ff$ be a category on $S$.
Denote by $\Ff^\alpha$  the category on $S$ having as sets of morphisms
$$\Hom_{\Ff^\alpha}(Q,R)=\alpha^{-1}\circ\Hom_\Ff(\alpha(Q),\alpha(R))\circ\alpha\,.$$
Note that if $\Ff\in\Cc_J$ then $\Ff^\alpha\in\Cc_J$, and if $\psi\in\Hom_\Ff(Q,R)$ then $\alpha \psi \alpha^{-1} \in \Hom_{\Ff^{\alpha^{-1}}}(\alpha(Q),\alpha(R))$.
Thus we have:
$$\begin{array}{lll}
\alpha(W)&:=\langle \alpha\psi(W_0) \; | \; \psi \in \Hom_\Ff(J(S),S) \;{\rm for}\;\Ff \in \Cc_J \rangle = &\\
&= \langle \alpha\psi\alpha^{-1}(\alpha(W_0)) \; | \; \alpha\psi\alpha^{-1} \in \Ff^{\alpha^{-1}}(\alpha(J(S)),\alpha(S)) \;{\rm for}\;\Ff \in \Cc_J \rangle = &\\
&= \langle \widetilde\psi(W_0) \; | \; \widetilde\psi \in \Ff^{\alpha^{-1}}(J(S),S) \;{\rm for}\;\Ff \in \Cc_J \rangle &\\
&\leq W
\end{array}
$$
where in the last equality we use that $W_0$ and $J(S)$ are characteristic subgroups of $S$.
But since $\alpha$ is injective, it follows that $|W| =|\alpha(W)|$ and therefore $\alpha(W)=W$, proving that $W$ is characteristic in $S$.
\end{proof}

\subsection{Stellmacher functor} Given that $(\varphi, \Ff)$ and $(\alpha\varphi, \Ff^{\alpha^{-1}})$ are equivalent as embeddings, $\Ff\in\Cc_J$ if and only if $\Ff^\alpha\in\Cc_J$ for any $\varphi,\alpha\in\Aut(S)$, and $\Ff$ a category on $S$, the two definitions $W(S)$ and $W$ represent the same subgroup of $S$, that is $W=W(S)$. It follows from Lemma $4.2$ that the functor $S \rightarrow W(S)$, for $S$ a finite $p$-group, is a positive characteristic functor in the sense of Definition \ref{charfunct}. We shall call the functor $S \rightarrow W(S)$, with $W(S)$ constructed via one of the methods from this Section, a {\it Stellmacher functor}.

The Thompson subgroup of $S$, $J(S)$ is a characteristic, centric subgroup. Thus using $(C1)$, any $\Ff\in\Cc_J$ is a constrained fusion system on $S$ and by Theorem \ref{thmconstraint}, there exists a $p$-constrained finite group $L$ with $\Ff = \Ff_L(S)$ and satisfying the following conditions: $S$ is a Sylow $p$-subgroup of $L$; $\C L{O_p(L)} \leq O_p(L)$ and $L$ is $Qd(p)$-free. It follows from our construction that $W(S)$ is a characteristic subgroup of $S$ which is also normal in $L$. The construction of $W(S)$ depends on $S$ only, and the subgroup $W(S)$ is constructed in the same way  as Stellmacher does in the context of finite groups so it is the same subgroup of $S$. Finally, notice that $S \rightarrow W(S)$ is also a Glauberman functor, see Definition \ref{charfunct}.

%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
\section{Proofs of the Theorems}%%%%%%%%%%%%%%%%%%%
%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%

\subsection*{Proof of Theorem 1.1}
Let $S$ be a finite $2$-group and let $\Ff$ be a $\Sigma_4$-free fusion system on $S$. Also $W(S)$ is the characteristic subgroup of $S$ defined in Section $\ref{charsg}$. We prove that the normalizer $\N\Ff{W(S)}$ of 
$W(S)$ in $\Ff$ is equal to $\Ff$.

This is true for the smallest fusion system on $S$, which is $\Ff_S(S)$. Suppose now by induction that all proper $\Sigma_4$-free subsystems and all $\Sigma_4$-free quotient systems $\Ff/Q$, with $Q$ a nontrivial normal subgroup of $\Ff$ satisfy Theorem \ref{mainS4}.

If $O_2(\Ff)=1$ then for every non-trivial fully $\Ff$-normalized subgroup $P$ of $S$ we have that $\N\Ff P$ is a proper subsystem of $\Ff$ (otherwise $P\le O_2(\Ff)$). But then $\N\Ff P$ satisfies Theorem \ref{mainS4} by induction as it is $\Sigma_4$-free by Proposition \ref{subsyst}. Hence $\N\Ff P=\N{\N\Ff P}{W(\N SP)}$ for every non-trivial fully $\Ff$-normalized subgroup $P$. An application of Proposition \ref{normal} gives now that $\Ff=\N\Ff{W(S)}$. So we can suppose $\bf(H1)$: $O_2(\Ff)\ne 1$.

Set $Q:=O_2(\Ff)$ and $R:=Q\C SQ$. These subgroups of $S$ are both nontrivial by $(H1)$.
If $Q=R$ then $Q$ is $\Ff$-centric. Consequently $\Ff$ is a constrained fusion system. According to \cite[Proposition 4.3]{bcglo} there exists a $2'$-reduced $2$-constrained finite group $L_Q$, which is an extension of $\Aut_\Ff(Q)$ by $Z(Q)$ and having $S$ as a Sylow $2$-subgroup. Thus $\Ff=\Ff_S(L_Q)$. Since $\Ff$ is $\Sigma_4$-free, the group $L_Q$ is $\Sigma_4$-free, by definition and given the $L_Q$ is $2$-constrained we have
$C_{L_Q}(O_2(L_Q))\le O_2(L_Q)$. Then, according to Stellmacher's main theorem in \cite{st96}, see also Section 1, the group $W(S)$ is normal in $L_Q$. This in its turn implies that $W(S)\trianglelefteq \Ff_S(L_Q)$ and therefore $\Ff=\N\Ff{W(S)}$. Thus we can also make the assumption $\bf(H2)$: $Q\ne R$
implying moreover that $\N\Ff R$ is a proper subsystem of $\Ff$.

Next we will see that we also have $\bf(H3)$: $S\C\Ff Q \ne \Ff$. Indeed suppose that $S\C\Ff Q=\Ff$. Then $\Ff/Q$ is a proper quotient system of $\Ff$ which is $\Sigma_4$-free by Proposition \ref{quotient}. The induction hypothesis gives now $\Ff/Q=\N{\Ff/Q}{W(S/Q)}$. Next, Proposition \ref{quot} gives that $\Ff=\N\Ff U$ where $U$ is the preimage in $S$ of $W(S/Q)$. As $U \trianglelefteq \Ff$ it follows that $U\le Q$, but this leads to a contradiction given that $W(S/Q)\ne 1$.

According to Lemma \ref{Op_cr}, we have $\Ff = \langle \Ff_1, \Ff_2 \rangle$ with $\Ff_1=S \C \Ff Q$ and $\Ff_2 = \N \Ff R$. By $(H2)$ and $(H3)$ both $\Ff_1$ and $\Ff_2$ are proper subsystems of $\Ff$, the induction hypothesis gives that $W(S) \trianglelefteq \Ff_1$ and that $W(S) \trianglelefteq \Ff_2$.

Notice that $W(S)$ is fully $\Ff$-normalized. By Lemma \ref{fullynormalized}, there exists a morphism $\varphi\in\Hom_\Ff(N_S(W(S)),S)$ with $\varphi(W(S))$ fully $\Ff$-normalized. As $N_S(W(S))=S$ and since $W(S)$ is characteristic in $S$ it follows that $W(S)=\varphi(W(S))$ is fully $\Ff$-normalized.

Finally, an application of Lemma \ref{normalizerW} gives the result: $\Ff=\N\Ff{W(S)}$.

\subsection*{Proof of Theorem 1.2}
Let $S$ be a finite $p$-group. Recall that the construction $W(S)$ described in Section $4$ and which associates to $S$ a nontrivial characteristic subgroup $W(S)$ gives rise to a Glauberman functor.

Assume now that $\Ff$ is a $Qd(p)$-free fusion system on $S$. If $p$ is an odd prime, it follows from \cite[Theorem B]{kl05} that $\Ff = \N\Ff{W(S)}$. If $p=2$ then $Qd(2)=\Sigma_4$ and the result is given by Theorem \ref{mainS4}.

\subsection*{Proof of Theorem 1.3}
Let $p$ be an odd prime. Let $\Ff$ be a fusion system over a finite $p$-group $S$. Let $W(S)$ be the characteristic subgroup of $S$ given by the Stellmacher functor. Since $\Ff_S(S)\subseteq \N \Ff {W(S)} \subseteq \Ff$ it is enough to show that if $\N \Ff{W(S)}=\Ff_S(S)$ then $\Ff = \Ff_S(S)$. The proof is similar to that of Theorem $A$ in \cite{kl05}; for the sake of completeness we will provide the details.

Let $\Ff$ be a minimal counterexample to Theorem $1.3$; thus $\N \Ff {W(S)}=\Ff_S(S)$ but $\Ff \ne \Ff_S(S)$, and all the proper subsystems and quotient systems of $\Ff$ satisfy Theorem $1.3$. Under this assumption we show that $\Ff$ is a constrained fusion system by proving that $Q :=O_p(\Ff)$ is a nontrivial $\Ff$-centric proper subgroup of $S$. This is attained in the following six steps.

{\it Step 1 : Any fusion system $\Gg$ on $S$ which is properly contained in $\Ff$ is equal to $\Ff_S(S)$}.

As $\Gg \subset \Ff$ it follows that $\N \Gg {W(S)} \subseteq \N \Ff {W(S)} = \Ff_S(S)$. But $W(S)$ is a characteristic subgroup of $S$ and therefore $\Ff_S(S) \subseteq \N \Gg {W(S)}$. Thus $\N \Gg {W(S)} = \Ff_S(S)$ and the minimality assumption on $\Ff$ implies that $\Gg = \Ff_S(S)$.

{\it Step 2 : Let $P$ be a fully $\Ff$-normalized subgroup of $S$ and set $A=N_S(P)$. Then there is $\varphi \in \Hom_\Ff(A,S)$ such that both $\varphi(P)$ and $\varphi(W(A))$ are fully $\Ff$-normalized}.

By Lemma $2.2(a)$, there is a morphism $\varphi:N_S(W(A)) \rightarrow S$ such that $\varphi(W(A))$ is fully $\Ff$-normalized. Since $W(A)$ is a characteristic subgroup of $A$, we have $N_S(P)=A \leq N_S(A) \leq N_S(W(A))$ and the morphism $\varphi$ can be restricted to $\varphi:A \rightarrow S$. According to Lemma $2.2(b)$, the group $\varphi(P)$ is also fully $\Ff$-normalized.

{\it Step 3 : The subgroup $Q = O_p(\Ff)$ is nontrivial}.

Recall that it is assumed that $\Ff_S(S) \subset \Ff$. Alperin's fusion theorem implies that there is a fully $\Ff$-normalized subgroup $P$ of $S$ with $\Ff_A(A) \subset \N \Ff P$, for $A = N_S(P)$. Choose the subgroup $P$ such that:\\
\hspace*{.4cm}a) $W(A)$ is fully $\Ff$-normalized;\\
\hspace*{.4cm}b) $N_S(P)=A$ has maximal order among subgroups $T$ with $\Ff_{N_S(T)}(N_S(T)) \subset \N \Ff T$.\\
The choice of $P$ and the fact that $A$ is a proper subgroup of $N_S(W(A))$ implies that $\N \Ff {W(A)} = \Ff_{N_S(W(A))}(N_S(W(A))$. Therefore $N_{\N \Ff P}(W(A))=\Ff_A(A)$.

If $\N \Ff P \subset \Ff$ then the minimality assumption on $\Ff$ implies that $\N \Ff P = \Ff _R(R)$, which contradicts our choice of $P$. Thus we have $\N \Ff P = \Ff$ and $1 \ne P \trianglelefteq \Ff$. Hence $1 \ne P \leq Q$ which proves that $Q \ne 1$.

{\it Step 4 : $Q$ is a proper subgroup of $S$}.

If $Q=S$ then $\Ff = \N \Ff S=\N \Ff {W(S)} = \Ff_S(S)$ contradicting our assumption on $\Ff$.

{\it Step 5 : $S \C\Ff Q = \Ff_S(S)$ when $Q = O_p(\Ff)$}.

We have $S \C\Ff Q \subseteq \Ff$. If $S \C\Ff Q \subset \Ff$ then {\it Step 1} implies that $S \C\Ff Q = \Ff_S(S)$ and we are done. Assume now that $S \C \Ff Q = \Ff$ and recall that $\Ff \ne \Ff_S(S)$. An application of Proposition \ref{quot}, with $\Gg=\Ff_S(S)$ and $\Ff=\N\Ff Q$, gives that $\Ff /Q \ne \Ff_S(S)/Q = \Ff_{S/Q} (S/Q)$. By {\it Step 3} the subgroup $Q$ is nontrivial and the minimality assumption on $\Ff$ implies that $\N {\Ff/Q} {W(S/Q)} \ne \Ff_{S/Q}(S/Q)$. Let $P$ be the inverse image of $W(S/Q)$ in $S$. Notice that $W(S/Q) \ne 1$, by the definition of $W(S)$, and thus $P$ properly contains $Q$. Also $P \trianglelefteq S$ and $\N {\Ff/Q} {W(S/Q)} = \N {\Ff/Q} {P/Q}$. Another application of Proposition \ref{quot} gives that $\N \Ff P \ne \Ff_S(S)$. Since $\N \Ff P \subseteq \Ff$, {\it Step 1} implies that $\N \Ff P = \Ff$ which is a contradiction with the fact that $P$ contains $Q$ properly.

{\it Step 6 : The subgroup $Q$ is $\Ff$-centric}.

If $Q = R=SC_S(Q)$ then $Q$ is $\Ff$-centric and we are done. So let us assume that $Q <R$. Notice that $R \trianglelefteq S$. Then $\N \Ff R$ is a proper subsystem of $\Ff$ and an application of {\it Step 1} gives that $\N \Ff R = \Ff_S (S)$. Recall also that by the previous step, $S \C \Ff Q = \Ff_S(S)$. Therefore, Lemma \ref{Op_cr} implies that $\Ff = \Ff_S(S)$, which is a contradiction to our choice of $\Ff$. Thus we must have $Q = R$.

Since $Q = O_p(\Ff)$ is a nontrivial normal centric subgroup of $\Ff$, the fusion system $\Ff$ is constrained. But this means by \cite[4.3]{bcglo1}, that there is a $p'$-reduced $p$-constrained finite group $L$ with $S$ as a Sylow $p$-subgroup and such that $Q=O_p(L)$. Furthermore $\Ff = \Ff_S(L)$ and therefore $\N \Ff {W(S)} = \Ff_S (\N L{W(S)}$.

Since $\N \Ff{W(S)}=\Ff_S(S)$ it follows that $\N L{W(S)}$ has a $p$-complement; see Remark $6.10$ in Appendix. According to the normal $p$-complement theorem of Thompson, $6.11$ below,  it follows that $L$ has a $p$-complement. Therefore $\Ff_S(S)=\Ff_S(L)$ and we reached a contradiction with our assumption on $\Ff$. This concludes the proof of the Theorem $1.3$.

%%%%%%%%%%%%%%%%%%%%
\section{Appendix}%%
%%%%%%%%%%%%%%%%%%%%

Let $p$ be an odd prime, $G$ a finite group and $S$ a Sylow $p$-subgroup of $G$. We say that $G$ is $p$-{\it stable} if and only if for every $p$-subgroup $Q$ of $G$ and every element $x$ of $\N GQ$ such that $[Q,x,x]=1$, we have that $x\C GQ \in O_p(\N GQ/\C GQ)$.

A classic result of special significance to the theory of finite groups is Glauberman's ZJ-theorem \cite{gl68}:

\begin{thma}[Glauberman] Let $p$ be an odd prime. Let $G$ be a finite, $p$-stable group such that $\C G{O_p(G)} \leq O_p(G)$. Then $Z(J(S))$ is a normal subgroup of $G$.
\end{thma}

Using the following:

\begin{propa}[14.7, \cite{gl71}] Assume that $p$ is odd and that $G$ is a finite group. Then the following conditions on $G$ are equivalent:\\
\hspace*{.5cm}(a) the group $Qd(p)$ is not involved in $G$;\\
\hspace*{.5cm}(b) every section of $G$ is $p$-stable.
\end{propa}

the $ZJ$-theorem can be reformulated as follows:

\begin{thma}[Glauberman] Let $p$ be an odd prime and let $G$ be a $Qd(p)$-free finite group with $\C G{O_p(G)} \leq O_p(G)$. Then $Z(J(S))$ is a normal subgroup of $G$.
\end{thma}

For $p=2$ the $ZJ$-theorem does not hold anymore; see \cite[Section 11]{gl71}. As noted by Glauberman \cite{gl71} a necessary and sufficient condition for every section of $G$ to be $2$-stable is that $G$ have a normal $2$-complement, which is too strong to be useful.

In a couple of papers \cite{st92,st96}, Stellmacher proved an analogous version of Glauberman's $ZJ$-theorem, by constructing a characteristic subgroup $W(S)$ of $S$ and extending the result for $p=2$. An overview of his method, including a sketch of the proof for the odd prime case can be found in \cite[Section 9.4]{ks04}. The main theorem in \cite{st96} reads as follows:

\begin{thma}[Stellmacher] Let $S$ be a nontrivial finite $2$-group. Suppose that $G$ is a finite group satisfying the following:\\
\hspace*{.4cm}  (I) $G$ is $\Sigma_4$-free,\\
\hspace*{.4cm} (II) $S \in \Syl_2(G)$ and $C_G(O_2(G)) \leq O_2(G)$,\\
\hspace*{.4cm}(III) Every non-abelian simple section of $G$ is isomorphic to $Sz(2^{2n+1})$ or $PSL_2(3^{2n+1})$.
Then there exists a nontrivial characteristic subgroup $W(S)$ of $S$ which is normal in $G$.
\end{thma}

Next, consider a couple of useful lemmas:

\begin{lema}[Chp. II, Lemma 2.3, \cite{gl77}] The following conditions are equivalent:\\
\hspace*{.5cm}(a) $\Sigma_4$ is involved in $G$;\\
\hspace*{.5cm}(b) There exists a $2$-subgroup $Q$ of $G$ such that $\Sigma_3$ is involved in $\N G Q/\C GQ$.
\end{lema}

\begin{lema}[Chp. II, Corollary 7.3, \cite{gl77}] Let $G$ be a non-abelian simple group. The following are equivalent:\\
\hspace*{.5cm}(a) $G$ is $\Sigma_3$-free;\\
\hspace*{.5cm}(b) $G$ is isomorphic to $Sz(2^{2n+1})$ or $PSL(2,3^{2n+1})$.
\end{lema}

\begin{rem} Note that if $G$ is $\Sigma_3$-free then $G$ is $\Sigma_4$-free. A finite group $G$ with $\C G{O_2(G)} \leq O_2(G)$ is $\Sigma_4$-free if and only if $G/O_2(G)$ is $\Sigma_3$-free \cite{st96}.
\end{rem}

Using the previous two lemmas and remark, we can rephrase Stellmacher's Theorem $6.4$ as follows:

\begin{thma}[Stellmacher] Let $S$ be a finite nontrivial $2$-group. Then there exists a nontrivial characteristic subgroup $W(S)$ of $S$ which is normal in $G$, for every finite $\Sigma_4$-free group $G$ with $S$ a Sylow $2$-subgroup and $\C G{O_2(G)} \leq O_2(G)$.
\end{thma}

If $G=S O_{p'}(G)$, with $S$ a Sylow $p$-subgroup of $G$, we say that $G$ {\it has a normal p-complement}. A standard result due to Frobenius (see \cite[8.6]{gl71} for example) is given below:

\begin{thma}[Frobenius] The following conditions are equivalent for a finite group $G$ with Sylow $p$-subgroup $S$:\\
\hspace*{.4cm}(a) $G$ has a normal $p$-complement;\\
\hspace*{.4cm}(b) if $Q$ is a non-identity subgroup of $G$ then $\N G Q /\C G Q$ is a $p$-group;\\
\hspace*{.4cm}(c) if $Q$ is a non-identity $p$-subgroup of $G$ then $\N G Q$ has a normal $p$-complement;\\
\hspace*{.4cm}(d) if two elements of $S$ are conjugate in $G$, they are conjugate in $S$.
\end{thma}

\begin{rem} The equivalence (a) $\Leftrightarrow$ (d) in the above theorem, states that $G$ has a normal $p$-complement if and only if $S$ controls fusion in $G$. In the language of fusion systems, $S$ controls $G$ fusion if and only if $\Ff_S(S)=\Ff_S(G)$.
\end{rem}

For odd primes Frobenius' result was improved by a result of Thompson. We give below a version of Thompson's p-complement theorem which uses Stellmacher's characteristic subgroup $W(S)$:

\begin{thma}\cite[9.4.7]{ks04} Let $G$ be a group, $p$ an odd prime, and $S \in \Syl_p(G)$. Then $G$ has a normal $p$-complement if and only if $\N G{W(S)}$ has a normal $p$-complement.
\end{thma}

\end{document}